# Adversary Resilient Learned Bloom Filters

### v3

@@ Appendix B Pseudo-random Permutations
Definition B.3.
Let $F:{\{0,1\}}^{*}\times{\{0,1\}}^{*}\mapsto{\{0,1\}}^{*}$ be an efficient keyed permutation. $F$ is a pseudo-random permutation if for all probabilistic polynomial time distinguishers $D$, there exists a negiligible function $\mathtt{negl}$, such that
@@
Appendix C Hybrid Model - Additional Experiments
This section contains additional experiments under the Hybrid Adversarial Model setting described in Section 5. We analyze the Hybrid Model using $494$ lines of C code in the model/ directory of our open-source implementation [Bod24].
C.1 Varying Non-Adversarial True Negatives ($Q_{N}$)
@@
C.2 Adversary Strategies
In the main text, we explored the case where the adversary divides $\alpha N$ queries equally between queries that generate False Positives, and queries that generate False Negatives. We conduct an experiment here for all partitions of $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$ for $\alpha=0.2$ to see how the False Positive Rate of the Downtown Bodega Filter is impacted. We vary the fraction of $\alpha$ assigned to $\alpha_{P}$ from 00 to $1$. Here, 00 means the adversary spends their entire budget of $\alpha N$ queries on true negatives, and $1$ means the adversary spends their entire budget of $\alpha N$ queries on false positives. Our earlier experiments are setting this fraction to $0.5$, in this framework. The results are in Figure 8
C.3 Additional Datasets
We conduct experiments on two other common evaluation datasets used in prior work on the Learned Bloom Filter [SM24, VKMK21, DS20]. We include a brief description of these datasets, adapted from Sato and Matsui’s recent NeurIPS 2023 work on Fast Partitioned Learned Bloom Filters [SM24].
  • Malicious URLs Dataset [mal]: The URLs dataset contains $223,088$ malicious and $428,118$ benign URLs.
  • EMBER Dataset [AR18]: This dataset contains $300,000$ malicious and $400,000$ benign files.
@@ C.3 Additional Datasets
Appendix D Implementation
We implement both the Downtown Bodega Filter and the Secure Classical Bloom Filter in $441$ lines of Python 3 code, which can be found in the src/ directory of our open-source implementation [Bod24]. Similar to Naor Yogev [NE19], we implement pseudo-random permutations using AES with a block size of $128$ bits. We use CBC-mode. Our implementation makes use of the PyCryptoDome [PyC] library for these cryptographic primitives. Our implementation is modular and allows any classification model to be easily plugged in. We tested the correctness of our implementation on a broad range of common classification models including the Random Forest model, Gaussian Naive Bayes, the Gradient Boosting Classifier, Support Vector Machine based Classifiers, and Adaptive Boosting. We relied on implementations of these models provided by scikit-learn.
D.1 Experimental Setup
Similar to prior work [SM24, VKMK21], we tested our implementation on the Malicious URLs dataset [mal], which we describe in Appendix C. We unoformly randomly sample $20\%$ of the dataset to reduce our experimentation time. We use a set of $20$ features to train our Learning Models including: length of the URL, whether the URL contains an IP address, whether the URL uses a shortening service, whether the URL is “abnormal”, the digit count and letter count of the URL, and whether the URL contains special symbols such as @,?, * etc. This set of features for the Malicious URLs Dataset is common in open-source learning models, and similar features have been used by Vaidiya, Knorr, Mitzenmacher et al in prior work [VKMK21].
D.2 Methodology
We calculate the False Positive Rate of both the Secure Classical Bloom Filter as well as the Downtown Bodega Filter by uniformly randomly sampling URLs that are not malicious from Malicious URLs Dataset and counting how many of them are returned as false positive by both filters. We calculate the memory used by the learning model by using joblib to create a persistent version [Sci] of the model, and measuring the size of the persistent version of the model. After subtracting the memory used by the learning model from the memory budget, we divide the remaining memory budget equally between Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$.
D.3 Results
Figure 11 shows how how the False Positive Rate varies for both the Secure Classical Bloom Filter as well as the Downtown Bodega Filter as we modify the memory budget available. This figure uses the Gaussian Naive-Bayes Classifier as the Bloom Filter Learning Model. Figure 12 shows the same results using a Linear Support Vector Classifier as the BLoom Filter Learning Model. [^1] All the data used to plot these results can be found in src/results_plot.py in our code repository.
Figure 11: The FPR of the Downtown Bodega Filter using a Gaussian Naive-Bayes Classifier as its Bloom Filter Learning Model compared to the Secure Classical Bloom Filter with varying memory budget as evaluated on the Malicious URLs Dataset
Figure 12: The FPR of the Downtown Bodega Filter using a Linear Support Vector Classifier as its Bloom Filter Learning Model compared to the Secure Classical Bloom Filter with varying memory budget as evaluated on the Malicious URLs Dataset

### v4

@@ Appendix 0.A Pseudo-random Permutations
Definition 23
Let $F:{\{0,1\}}^{*}\times{\{0,1\}}^{*}\mapsto{\{0,1\}}^{*}$ be an efficient keyed permutation. $F$ is a pseudo-random permutation if for all probabilistic polynomial time distinguishers $D$, there exists a negligible function $\mathtt{negl}$, such that
@@
Appendix 0.B Hybrid Model - Additional Experiments
This section contains additional experiments under the Hybrid Adversarial Model setting described in Section 5. We analyze the Hybrid Model using $494$ lines of C which we have open-sourced [16].
0.B.1 Varying Non-Adversarial True Negatives ($Q_{N}$)
@@
0.B.2 Adversary Strategies
In the main text, we explored the case where the adversary divides $\alpha N$ queries equally between queries that generate False Positives and queries that generate False Negatives. We conduct an experiment here for all partitions of $\alpha$ between $\alpha_{P}$ and $\alpha_{N}$ for $\alpha=0.2$ to see how the False Positive Rate of the Downtown Bodega Filter is impacted. We vary the fraction of $\alpha$ assigned to $\alpha_{P}$ from 00 to $1$. Here, 00 means the adversary spends their entire budget of $\alpha N$ queries on true negatives, and $1$ means the adversary spends their entire budget of $\alpha N$ queries on false positives. Our earlier experiments are setting this fraction to $0.5$, in this framework. The results are in Figure 8
0.B.3 Additional Datasets
We conduct experiments on two other common evaluation datasets used in prior work on the Learned Bloom Filter [17, 18, 19]. We include a brief description of these datasets, adapted from Sato and Matsui’s recent NeurIPS 2023 work on Fast Partitioned Learned Bloom Filters [17].
  • Malicious URLs Dataset [20]: The URLs dataset contains $223,088$ malicious and $428,118$ benign URLs.
  • EMBER Dataset [21]: This dataset contains $300,000$ malicious and $400,000$ benign files.